# On-device Anomaly Detection in Conveyor Belt Operations

### v1

@@ IV Results and Discussion
TABLE II: Average F1-score for the duty cycle detection in each dataset. The detection does not depend on the operation mode classifier in Approach-2 and Approach-3.
@@ IV-A Recognition Performance
The detection and classification performance of the approaches using different ML models is presented in terms of F1-scores, averaged over repeated experiments with varying seeds of initialization. Table II shows the detection rates for operation cycles across approaches. In Approach-1, correct cycle detection relies on accurate classification of operation modes, yielding consistent performance above 89.3% across all ML models in DS1. In DS2, the detection performance varies across ML models. While RF, DT, and XGB models show improvement over DS1, MLP, ET, and NB experience declines. In contrast, Approach-2 and Approach-3 detect cycles independently of the operation mode classifier, using threshold-based detection (see Fig. 2(b) and Fig. 2(c)). This results in 2.5% to 8.2% improvements in DS1 and 1.3% to 37.6% in DS2 over Approach-1.
Figures 3 to 5 illustrate the duty cycle classification results for the three approaches in both datasets, revealing significant performance differences. DS1 consistently yields better results than DS2, reflecting the limited generalization capabilities of the approaches and suggesting distinct complexities and class separability challenges between the two datasets. Across all approaches, the abnormal cycle classification is consistently more challenging, with lower F1-scores than those for normal cycles, due to (i) class imbalance and (ii) the restrictive two possible operation mode sequence combinations in normal cycles compared to the diverse sequences in the abnormal class (see section III-B).
@@
TABLE III: Comparison between the selected approaches deployed on the MCUs in terms of classification performance, energy consumption, and memory usage.
@@
    Using MLP and ET as operation mode and duty cycle classifiers, respectively.
Approach-2 demonstrates substantial improvements over Approach-1 across both datasets (Fig. 3b). In DS1, Approach-2 maintains a comparable F1-score for normal cycles and shows a 7.5-24.6% improvement in abnormal cycle classification. ET again leads with an overall F1-score of 94.0%, achieving 97.3% for normal and 80.2% for abnormal cycles, improving overall performance by 1.7% and abnormal classification by 9.0% over Approach-1. In DS2, normal cycle classification with Approach-2 models is 0.5-4.1% lower than Approach-1. However, when using the MLP, ET, or NB models, the abnormal classification improves by 12.4%, 20.2%, and 35.6%, respectively, boosting overall F1-scores by 3.9%, 8.9%, and 15.0%. Approach-2 with the ET achieves the highest abnormal cycle F1-score, while RF excels in normal cycle and overall F1-scores.
For Approach-3, which combines separate classifiers for operation modes and duty cycles, the highest performance in DS1 is using ET for both classifiers, reaching an overall F1-score of 93.0% (96.6% for normal, 76.7% for abnormal) (Fig. 4). However, these results do not surpass the best performance obtained in Approach-2 using ET. Interestingly, in DS2, the combination of MLP for operation mode and ET for duty cycle classification yields the best performance in Approach-3, with an overall F1-score of 86.2% (91.3% for normal, 67.9% for abnormal) (Fig. 5). This result, the highest in DS2 across all approaches, suggests the dual-classifier structure may better address the complexity of DS2. These findings emphasize the importance of aligning classification methods with dataset characteristics and underscore the potential benefits of sophisticated architectures in complex classification scenarios.
@@ IV-B MCU Energy Requirements
The best-performing classifier-dependent approaches in each dataset are selected to be deployed in the MCUs using 32-float and 8-integer bit representations. These are Approach-2 using the ET as operation mode classifier and Approach-3 using the MLP and ET as operation mode and duty cycle classifiers, respectively. The effect of quantizing to 8-bit does not affect the duty cycle detection, keeping it at 99.2% as in the previous experiments (Table II). The selected models, trained in DS1 and deployed in the MCUs, are evaluated in DS2. The trade-off between classification performance and hardware resource utilization is presented in Table.
The best-performing classifier-dependent approaches for each dataset are selected for deployment on MCUs using either 32-bit float or 8-bit integer representation. These include Approach-2, which uses an ET classifier for operation mode classification, and Approach-3, which combines an MLP for operation mode classification and ET for duty cycle classification.
Quantizing to 8-bit does not impact duty cycle detection accuracy, which remained at 99.2% as observed in earlier experiments (Table II). Table III illustrates the balance between classification performance in DS2 and hardware resource usage. Approach-3 achieves superior classification rates for normal, abnormal, and overall classes compared to Approach-2, consistent with previous results. Interestingly, both approaches show improved classification performance when quantized to 8-bit, likely due to reduced quantization error. Energy consumption calculations focused on data representing duty cycles, ensuring that all system stages are active (Fig. 2), capturing the peak energy consumption. Reported energy consumption reflects the end-to-end computation for each 1-min input data, averaged across DS2. The time between inputs allows for additional tasks such as MCU peripheral management, conveyor control, or putting the MCU into sleep mode. The Arduino Nano BLE Sense 33 Lite exhibits lower energy consumption for both approaches in the 32-bit float version due to the MCU featuring a floating-point computation unit. In contrast, the Raspberry Pi Pico consumes significantly more power with 32-bit floats but outperforms the Arduino for 8-bit integer representations. Quantizing Approach-2 reduces program memory usage by 46.4% on the Arduino Nano BLE Sense 33 Lite and 44.4% on the Raspberry Pi Pico, while Approach-3 presents reductions of 5.1% and 4.8%, respectively. Across both MCUs, 32-bit float and 8-bit integer versions required similar amounts of RAM, as both store data in 32-bit registers.
@@ V Conclusion
This work contributes to Mining 4.0 by proposing two novel approaches for real-time anomaly detection in the continuous operation of hydraulic conveyor belt systems using MCUs. These pattern recognition methods combine feature extraction, threshold-based cycle detection, and TinyML classifiers to accurately identify normal and abnormal duty cycles. By evaluating the performance of different approaches and ML classifiers across two distinct datasets, we identified the most effective methods for achieving accurate and generalizable classification, essential for conveyor belt system monitoring and predictive maintenance. The first approach achieved higher performance when trained and tested on the same dataset, with an F1-score of 97.3% for normal and 80.2% for abnormal cycles, while the second approach showed superior performance on an independent dataset, achieving 91.3% for normal and 67.9% for abnormal cycles. This comparison highlights both the strengths and adaptability challenges of the methods across varying operational contexts. Furthermore, both approaches outperformed the state-of-the-art method on both datasets, validating their suitability for real-world scenarios.
@@
Taking advantage of the fact that the deployed approaches in MCUs use a small amount of the available resources, future work will aim to improve abnormal cycle classification by employing more complex models, such as recurrent neural networks or transformer-based models, to better capture temporal dependencies. Another research direction involves addressing the limited availability of abnormal cycle data through few-shot learning, as well as adding new unlabelled data into the model development using unsupervised ML techniques, which may enhance anomaly detection without extensive labeled datasets.

### v2

@@ IV Results and Discussion
TABLE IV: Average F1-score for the duty cycle detection in each dataset.
@@ IV-A Recognition Performance
The detection and classification performance of the approaches using different ML models is presented in terms of F1-scores, averaged over repeated experiments with varying seeds of initialization. Table IV shows the detection rates for operation cycles across approaches. In Approach-1, correct cycle detection relies on accurate classification of operation modes, yielding consistent performance above 89.3% across all ML models in DS1. In DS2, the detection performance varies across ML models. While RF, DT, and XGB models show improvement over DS1, MLP, ET, and NB experience declines. In contrast, Approach-2 and Approach-3 detect cycles independently of the operation mode classifier, using threshold-based detection (see Fig. 2(b) and Fig. 2(c)). This results in 2.5% to 8.2% improvements in DS1 and 1.3% to 37.6% in DS2 over Approach-1.
Figures 3 to 5 illustrate the duty cycle classification results for the three approaches in both datasets, revealing significant performance differences. DS1 consistently yields better results than DS2, reflecting the limited generalization capabilities of the approaches. This performance gap may stem from variations in the data range between the two datasets. These variations could be caused by differences in the transported materials, which may influence signal characteristics and separability. These findings suggest distinct complexities and class separability challenges between the two datasets. Across all approaches, abnormal cycle classification is consistently more challenging and yields lower F1-scores than normal cycles. This is due to (i) class imbalance and (ii) the restrictive two possible operation mode sequence combinations in normal cycles, compared to the diverse sequences in the abnormal class (see section III-B).
@@
Approach-2 demonstrates substantial improvements over Approach-1 across both datasets (Fig. 3b). In DS1, Approach-2 maintains a comparable F1-score for normal cycles and shows a 7.5-24.6% improvement in abnormal cycle classification. ET again leads with an overall F1-score of 94.0%, achieving 97.3% for normal and 80.2% for abnormal cycles, improving overall performance by 1.7% and abnormal classification by 9.0% over Approach-1. In DS2, normal cycle classification with Approach-2 models is 0.5-4.1% lower than Approach-1. However, abnormal classification improves by 12.4%, 20.2%, and 35.6% when using the MLP, ET, or NB models, respectively. This boosts the overall F1-scores by 3.9%, 8.9%, and 15.0%. Approach-2 with the ET achieves the highest abnormal cycle F1-score, while RF excels in normal cycle and overall F1-scores.
For Approach-3, which combines separate classifiers for operation modes and duty cycles, the highest performance in DS1 is using ET for both classifiers, reaching an overall F1-score of 93.0% (96.6% for normal, 76.7% for abnormal) (Fig. 4). However, these results do not surpass the best performance obtained in Approach-2 using ET. Interestingly, in DS2, the combination of MLP for operation mode and ET for duty cycle classification yields the best performance in Approach-3, with an overall F1-score of 86.2% (91.3% for normal, 67.9% for abnormal) (Fig. 5). This result, the highest in DS2 across all approaches, suggests the dual-classifier structure may better address the complexity of DS2. These findings emphasize the importance of aligning classification methods with dataset characteristics. They also highlight the potential benefits of sophisticated architectures in complex classification scenarios. The proposed methods are specific to a hydraulic conveyor belt. Their adaptability to new machines will depend on whether they allow monitoring of the same variables. TinyML models will need to be retrained on new machines with new configurations (physical dimensions or conveying speed).
TABLE V: Comparison between the selected approaches deployed on the MCUs.
@@
    Using MLP and ET as operation mode and duty cycle-classifiers, respectively.
@@ IV-B MCU Energy Requirements
The best-performing classifier-dependent approaches for each dataset are selected for deployment on MCUs using 32-bit float and 8-bit integer representation. These include Approach-2, which uses an ET classifier for operation mode classification, and Approach-3, which combines an MLP for operation mode classification and ET for duty cycle classification.
Quantizing to 8-bit does not impact duty cycle detection accuracy, which remained at 99.2% as observed in earlier experiments (Table IV). Table V illustrates the balance between classification performance in DS2 and hardware resource usage. Approach-3 achieves superior classification rates for normal, abnormal, and overall classes compared to Approach-2, consistent with previous results. Interestingly, both approaches show improved classification performance when quantized to 8-bit, likely due to reduced quantization error.
Energy consumption calculations focused on data representing duty cycles, ensuring that all system stages are active (Fig. 2), capturing the peak energy consumption. Reported energy consumption reflects the end-to-end computation for each 1-min input data, averaged across DS2. After completing the inference, the idle time can be utilized for additional tasks. These include MCU peripheral management, conveyor control, or putting the MCU into sleep mode. In the 32-bit float version, the Arduino Nano BLE Sense 33 Lite consumes less energy than the Raspberry Pi Pico. This is because it features a floating-point computation unit. In contrast, the Raspberry Pi Pico outperforms the Arduino for 8-bit integer representations. Quantizing Approach-2 reduces program memory usage by 46.4% on the Arduino Nano BLE Sense 33 Lite and 44.4% on the Raspberry Pi Pico, while Approach-3 presents reductions of 5.1% and 4.8%, respectively. Across both MCUs, 32-bit float and 8-bit integer versions required similar amounts of RAM, as both store data in 32-bit registers. We hypothesize that the power consumption of quantized approaches could be further reduced by using ultra-low-power 8-bit MCUs, such as the STM8L series MCUs [27].
@@ V Conclusion
This work contributes to Mining 4.0 by proposing two novel approaches for real-time anomaly detection in the continuous operation of hydraulic conveyor belt systems using MCUs. These pattern recognition methods combine manual feature extraction, threshold-based cycle detection, and supervised TinyML classifiers to accurately identify normal and abnormal duty cycles. By evaluating the performance of different approaches and ML classifiers across two distinct datasets, we identified the most effective methods for achieving accurate and generalizable classification. These methods are essential for conveyor belt system monitoring and predictive maintenance. The first approach performed better when trained and tested on the same dataset, achieving an F1-score of 97.3% for normal and 80.2% for abnormal cycles. In contrast, the second approach showed superior performance on an independent dataset, with F1-scores of 91.3% for normal and 67.9% for abnormal cycles. This comparison highlights both the strengths and adaptability challenges of the methods across varying operational contexts. Furthermore, both approaches outperformed the state-of-the-art method on both datasets, validating their suitability for real-world scenarios.
@@
Taking advantage of the fact that the deployed approaches in MCUs use a small amount of the available resources, future work will explore advanced feature engineering techniques to enhance discriminative power and robustness, thereby improving generalization across diverse operating conditions and datasets. Additionally, improving duty cycle classification will be investigated by employing more complex models, such as recurrent neural networks or transformer-based models, to capture temporal dependencies better. Another research direction involves addressing the limited availability of abnormal cycle data through few-shot learning, as well as adding new unlabeled data into the model development using unsupervised ML techniques, which may enhance anomaly detection without extensive labeled datasets.